OpenStudy (anonymous):

classify each number as rational interger whole natural or irrational 5,281 square root of 15 0 -5.2 2 5/7

OpenStudy (yanasidlinskiy):

First of all, Do you know what irrational, rational, and whole natural are?

OpenStudy (anonymous):

irrational can be simplified i tthink and rational can and whole can be simplified irdk

OpenStudy (yanasidlinskiy):

"Natural numbers"-1, 2, 3, 4, 5, 6, ... Rational Numbers- \[\frac{ 2 }{ 5 }, -\frac{ 7 }{ 9 },\frac{ 3 }{ 8 },2.3\] With a bar on top of the 3. Irrational Numbers- \[\sqrt{2}, \sqrt{5},\pi\]

OpenStudy (anonymous):

so 5281 is a natural number

OpenStudy (yanasidlinskiy):

Yep!!

OpenStudy (anonymous):

and is \[\sqrt{15}\] a irrational number

OpenStudy (yanasidlinskiy):

Yep!!! Correct, again!!

OpenStudy (anonymous):

and is 0 a whole

OpenStudy (yanasidlinskiy):

Look with what I gave you above!

OpenStudy (yanasidlinskiy):

That is your cheat sheet!:)

OpenStudy (anonymous):

so natural number

OpenStudy (yanasidlinskiy):

Yes.

OpenStudy (anonymous):

-5.2 is natural and 2 5/7 is rational

OpenStudy (yanasidlinskiy):

Yes. You shouldn't have any problems!:)

OpenStudy (anonymous):

thnk ur very smart algebra is hard

OpenStudy (yanasidlinskiy):

Your Welcome!!:) Glad to help!:)